OpenStudy (anonymous):

Please explain. thank you! medal for best answer! Is triangle ABC scalene, isosceles, or equilateral? The vertices are A(0 , 1), B(4 , 4), and C(7 , 0).

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@Callisto

OpenStudy (anonymous):

@undeadknight26

OpenStudy (phi):

did you plot the points ?

OpenStudy (anonymous):

no- i didnt kno if there was an easier way to do this

OpenStudy (anonymous):

i dont have any graph paper either

OpenStudy (phi):

just a sketch to see what it looks like. If it looks (roughly) like some of the sides might be equal, check using the distance formula

OpenStudy (anonymous):

it looks like a diagonal line

OpenStudy (phi):

If you plot the 3 points, you will get a triangle.

OpenStudy (anonymous):

oh i see wat i did wrong- so how do u tell which it is?

OpenStudy (anonymous):

im guessing isosceles

OpenStudy (anonymous):

its definitly no equilateral

OpenStudy (anonymous):

it looks like two sides are equal though

OpenStudy (phi):

Here is what it looks like It is clear (to me) that side AB is shorter than AC so yes, not equilateral. but AB and BC look suspiciously close in length. we could check by using the distance formula

OpenStudy (anonymous):

yeah- two sides are equal- that means its isosceles right?

OpenStudy (phi):

I'll do side BC and you can do side AB B is at (4,4) C is at (7,0) \[ \sqrt{ (x_2-x_1)^2 + (y_2-y_1)^2 }\\ \sqrt{ (7-4)^2 + (0-4)^2 }\\ \sqrt{9+16} = \sqrt{25}= 5 \]

OpenStudy (phi):

yes, triangle ABC is isosceles

OpenStudy (anonymous):

can i tag u in another question?

OpenStudy (phi):

yes